Decide whether or not the ordered pair satisfies the system of linear equations:
Yes or No
1. (2,-3) 2. (1/3,1/2)
{x=y+-1 {2x=10-y
{x-y+5 {x-4y=-5

Answer by Andr3w128(2)

1. (2,-3)
{x=y+-1
{x-y+5
(x=y+-1)
(2)=(-3+-1)
(2)=(-4)
DOES NOT SATISFY
(x=y+5)
(2)=(-3)+5
(2)=(2)
DOES SATISFY


2. (1/3,1/2)
(2x=10-y)
2(1/3)=10-(1/2)
(2/3)=(19/2)
DOES NOT SATISFY
(x-4y=-5)
((1/3)-4(1/2))=-5
(-5/3)=(-5)
DOES NOT SATISFY


-I like math